# RiskLabs: Predicting Financial Risk Using Large Language Model Based on Multi-Sources Data

### v1

@@ Abstract
The integration of Artificial Intelligence (AI) techniques, particularly large language models (LLMs), in finance has garnered increasing academic attention. Despite progress, existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&\&&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for financial risk prediction. Addressing this gap, in this paper, we introduce RiskLabs, a novel framework that leverages LLMs to analyze and predict financial risks. RiskLabs uniquely combines different types of financial data, including textual and vocal information from Earnings Conference Calls (ECCs), market-related time series data, and contextual news data surrounding ECC release dates. Our approach involves a multi-stage process: initially extracting and analyzing ECC data using LLMs, followed by gathering and processing time-series data before the ECC dates to model and understand risk over different timeframes. Using multimodal fusion techniques, RiskLabs amalgamates these varied data features for comprehensive multi-task financial risk prediction. Empirical experiment results demonstrate RiskLab’s effectiveness in forecasting both volatility and variance in financial markets. Through comparative experiments, we demonstrate how different data sources contribute to financial risk assessment and discuss the critical role of LLMs in this context. Our findings not only contribute to the AI in finance application but also open new avenues for applying LLMs in financial risk assessment.
1 Introduction
The integration of artificial intelligence (AI) techniques have dramatically transformed financial activities, including the way transactions are processed, investment decisions are made, and services are delivered across the industry Belhaj and Hachaıchi (2021); Ashta and Herrmann (2021). This technological evolution has facilitated the automation of complex and time-consuming tasks, allowing for more efficient and accurate operations.
Several compelling examples showcase how AI applications in finance support market participants to meet their needs for optimizing investment strategies, improving risk management, and enhancing customer service experiences. Investors are focused on identifying the underlying factors that influence key financial metrics. Techniques such as Random Forest, Support Vector Machines (SVM), and k_Nearest Neighbors (KNN) are effective at capturing both of the linear, non-linear relationships between variables Wu et al. (2006); Breiman (2001); Cortes and Vapnik (1995)Gu et al. (2020) explore the correlation between economic indicators and the performance of different asset classes using Tree-based models. Khaidem et al. (2016); Lee (2007) use SVM to predict credit rating using ratios from financial statements.
In addition to structured data like historical stock prices, investors are also seeking to derive trading signals from unstructured sources, such as financial reports and media news. Souma et al. (2019); Mohan et al. (2019); Liapis et al. (2023) use Recurrent Neural Networks (RNN) to analyze and extract sentiment from news articles, aiming to uncover market attitude towards the stock trends. Ahbali et al. (2022); Wang et al. (2023) use Natural Language Processing (NLP) techniques to monitor corporate credit risk by analyzing vast quantities of textual data for early warning signals and indicators of potential risks. Aue et al. (2022) use information extracted from news articles to construct Environmental, Social, and Governance (ESG) scores for companies. For instance, during quarterly earning conference calls, CEOs present the company’s past performance and future outlook, offering a wealth of information not just in the content but in the delivery itself. Qin and Yang (2019) focus on extracting information from audio recordings, such as changes in tone, speaker emotion, and the rate of speech, to predict the stock movement. They have shown that integrating audio analysis with traditional transcript text significantly enhances the accuracy of predictive models. Yang et al. (2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance.
The previously discussed models all fall under the category of supervised learning, which, while effective, is highly task-specific and has limited adaptability to the generalization Singh et al. (2016). For example, a supervised learning model trained to summarize financial documents would require extensive training on specialized, labeled data from that domain. Its performance could significantly deteriorate if applied to summarizing energy sector reports, illustrating a key limitation in domain adaptability. Furthermore, the predictive performance of these models is constrained by the relatively limited amount of input data length they can process and the size of their parameters Li et al. (2020). However, the emergence of Large Language Models (LLMs) represents a paradigm shift in overcoming these challenges. LLMs, with their expansive knowledge base and advanced zero-shot learning capabilities, can effectively perform a broad spectrum of text-related tasks, such as summarization Zhang et al. (2024), question-answer Wei et al. (2022), and sentiment analysis Zhang et al. (2023) without the necessity for task-specific or domain-centric training.
There has been a lot of prior work that has begun to explore how LLM can be used to solve problems related to the financial sector.  Li et al. (2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, 10-Q filings. Furthermore, Yang et al. (2023); Yu et al. (2023) explores the usage of LLMs in mining extensive media news to derive trading recommendations, highlighting the models’ ability to understand subtle market indicators and sentiments that might reveal human analysis. In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants Lakhani (2023); Subagja et al. (2023); Soni (2023)Abdaljalil and Bouamor (2021); Zmandar et al. (2021) have delved into the specific task of navigating financial and legal items from long text documents, such as financial regulations, legal stipulations, and comprehensive policy manuals. This research highlights the adeptness of LLMs in parsing and understanding complex textual information to identify and retrieve highly specific content. In summary, these scholarly works illuminate the diverse and impact applications of LLMs, affirming the critical role of advanced AI technologies in enhancing business operations and decision-making processes.
While there is a growing body of literature on the application of Large Language Models (LLMs) in finance, the specific area of financial risk prediction—such as market volatility and Value at Risk (VaR)—remains relatively unexplored. This observation underscores the novelty and significance of our research focus, which is dedicated to leveraging LLMs for financial risk prediction. The exploration of financial risk is crucial not only for understanding market dynamics but also for building resilience and ensuring sustainable success across various sectors Mitchell (2013); Van Der Vegt et al. (2015); Becker (2023). Our work seeks to bridge this research gap, recognizing the potential of LLMs to offer innovative insights in this pivotal aspect of finance.
In contemplating the application of Large Language Models (LLMs) for financial risk prediction, we recognized several inherent challenges. Primarily, analyzing market reactions like predicting stock volatility necessitates a comprehensive understanding of the multifaceted environment in which these markets operate, due to their inherent complexity. This requires the assimilation of diverse data sources, as different types of information can shed light on various aspects of the market. The true challenge lies in integrating multiple data types—ranging from quantitative data like stock prices and financial indicators to qualitative data from news articles, earnings transcripts, and social media—to form a more holistic view of the market. However, there are relatively few studies that attempt to merge these disparate data sources into a unified analysis. Bridging this gap could significantly enhance our ability to understand and predict market dynamics.
Moreover, the innate limitations of LLMs pose further hurdles in their functional deployment. For instance, For instance, using LLMs merely as tools for querying financial forecasts, with questions like “What will the stock price be tomorrow?", could yield imprecise outcomes. Utilizing large language models without careful data preprocessing and critical analysis of the results can lead to issues, leading users to make uninformed or misinformed financial decisions. One significant risk associated with this approach is the phenomenon known as ‘hallucination’, where LLMs generate responses incorrect responses rather than using factual accuracy or real-world data analysis Rawte et al. (2023).
Except for the hallucination problem, another inherent limitation when using Large Language Models to address financial queries lies in their inability to analyze current events and market conditions, as the training data for these models typically concludes at a specific cutoff point, such as 2021 Achiam et al. (2023). This presents a significant challenge in the context of financial analysis, where the relevancy and accuracy of information are significant, and the value of data diminishes rapidly with time. Financial markets are highly dynamic, with stock prices, economic indicators, and market sentiment changing constantly, often influenced by recent events, announcements, and global trends.
Given this gap, in designing financial risk prediction framework using LLM, this paper is interested in exploring the following research questions (RQs):
  • RQ1: How does the predictive performance of Large Language Models (LLMs) compare with other artificial intelligence techniques in forecasting risk metrics?
  • RQ2: What are the differences in predictive performance between traditional finance methods and Large Language Models LLMs in forecasting financial risk metrics?
  • RQ3: What role do large language models play in the financial area?
  • RQ4: How can multiple inputs of varying data types be effectively incorporated and balanced?
  • RQ5: How can the LLM model be updated with new media news and information?
In order to address the problems we mentioned above, this paper proposes a RiskLabs (Risk-Large Language model Based with Multi-sources input) framework. To accurately grasp the comprehensive complexities of the investment environment, our model is designed to ingest and analyze data from multiple sources, ensuring a holistic view of the factors that impact financial markets. This multifaceted data source includes: (1) Earnings conference call transcripts, which provide insight into company performance and future outlooks as communicated by corporate executives; (2) Earnings conference call audio, offering additional details such as tone and sentiment that might not be captured in text alone; (3) media news, to gauge the current economic, political, and social climate affecting market sentiment and investor behavior; and (4) time series data, for historical and real-time analysis of market trends and stock performance.
RiskLabs is equipped with four key modules to effectively process these diverse data streams: a) The Earnings Conference Call Encoder: leveraging LLM to handle data related to earnings conference call (1 and 2); b) The News-Market Reactions Encoder: establishing a pipeline through LLM to collect and interpret news data; c) The Time-Series Encoder: organizing and analyzing time-related data; and d) The Multi-Task Prediction: which amalgamates outputs from the aforementioned modules for multifaceted prediction. This synthesis of varied inputs allows the model to provide a nuanced understanding of the investment landscape, blending quantitative market data with qualitative insights. Such an integrated methodology is crucial for precise forecasting and risk assessment, a key to navigating the often turbulent and complex terrain of financial markets.
In our various modules, we employ Large Language Models (LLMs) to create dedicated pipelines for text analysis. To effectively process and analyze text-based content, we propose three pipelines: (1)Earnings conference call transcript analyzer, (2) enrich news attribute pipeline (3) news analyzer. These pipelines are based on an LLM technique known as retrieval augmentation generation (RAG) Lewis et al. (2020); Shuster et al. (2021). This approach is specifically designed to mitigate the issue of hallucination and use the updated material. More specifically, our approach goes beyond simply inputting a text file into an LLM model and awaiting the output. Instead, it is equipped with a function designed to interact intelligently with the LLM, thereby refining the retrieval process. This interaction enables the system to dynamically adjust queries based on initial responses, ask follow-up questions, or even reformulate the input to ensure that the information being sought is accurately retrieved and analyzed.
The news analyzer plays a crucial role in identifying the connections between news articles and market reactions, adopting a dual perspective of recency and relevance to achieve this. For recency, the focus is on analyzing news from the last three days, ensuring that the most current events are considered for their potential impact on the market. Regarding relevance, the goal is to identify news articles from the past that bear the highest similarity to current events. Similar events tend to have similar market responses. However, this task is challenging due to the vast diversity of news, making it difficult to find historical counterparts based solely on similarity. Therefore, to tackle this, the analyzer leverages an enriched news attributes pipeline, which encompasses a wide array of information related to each news piece. This information acts as a preliminary filter, helping to eliminate unrelated or irrelevant articles early in the process. Subsequently, the top ’k’ news articles that most closely match the current news in question are selected. This methodical approach allows for a refined analysis that significantly reduces noise and enhances the accuracy of understanding how current events resonate with historical market reactions, offering insightful perspectives on potential future market movements.
In addition to processing text data, our system includes a specialized module designed to handle audio data through the use of the LLM model. This innovative module extends the capabilities of traditional LLMs beyond textual analysis, allowing for the interpretation and understanding of earnings conference call recordings.
By incorporating a diverse range of inputs and using processing steps, RiskLabs solves a multitasking problem: forecasting volatility over various intervals (3 days, 7 days, 15 days, and 30 days), as well as the value at risk (VaR). To enhance the predictive capabilities for solving multitasking problems, we have developed a novel approach dubbed the Bayes-VaR method. This innovative technique is specifically designed to capture and elucidate the intricate relationships between response variables, which are crucial for forecasting tasks such as volatility and Value at Risk (VaR) estimations over different time intervals. By leveraging Bayesian statistics, the Bayes-VaR method provides a probabilistic framework that allows for the incorporation of prior knowledge and the updating of beliefs as new evidence is presented. This approach not only improves the accuracy of predictions by taking into account the inherent uncertainties and correlations in financial markets but also offers a more nuanced understanding of the risk factors influencing these metrics. Through this method, we can better model the complex dynamics between various financial indicators, thereby significantly enhancing our ability to predict market behaviors and risk exposures with greater precision and reliability.
Our contributions are encapsulated as follows: 1) We utilize LLM to develop the RiskLabs framework, thereby addressing the shortfall in applying LLMs to financial risk prediction. 2) The versatility of our model lies in its capability to seamlessly integrate a variety of financial data from multiple sources, consequently bolstering risk prediction accuracy. 3) The efficacy of our framework is evidenced by experimental results, demonstrating its effectiveness in forecasting financial risks. 4) Furthermore, we provide a comprehensive analysis of the pivotal role LLM should assume in the realm of financial risk prediction.
The subsequent sections of the article are outlined as follows: Section 2 provides an in-depth exploration of the data types and the specific dataset utilized in our study. It also clarifies the nature of our prediction task and delves into the intricate details of our framework’s structure and functionality. In Section 3, we conduct a thorough evaluation of both the overall performance of the model and the individual efficacy of each module. This section also includes comprehensive responses to the research questions outlined earlier. Finally, Section 4 revisits the central themes of our research, summarizing key insights and proposing practical solutions to address the current challenges and obstacles in the field.
2 RiskLabs Framework
Figure 1 illustrates the RiskLabs Framework, designed to handle multiple data types surrounding the financial information including audio, text, and time-series from different sources. The framework comprises four main modules: 1) Earnings Conference Call Encoder; 2) Time-Series Encoder; 3) Relevant News Encoder; and 4) Multi-Task Prediction Block. This section outlines the format of the data input and details each module. In brief, the Earnings Conference Call Encoder, Time-Series Data Encoder, and Relevant News Encoder are utilized to extract features from various data types. These features are fused data undergoes processing and modeling, after which it is fed into the Multi-Task Prediction Block. This block is responsible for forecasting both volatility for difference interval and VaR (Value at Risk) values.
@@ 2 RiskLabs Framework
Figure 1: This figure illustrates the RiskLabs Framework. The model accepts multiple-source inputs: Earnings Conference Call Audio & Transcript, Daily News, and Time Series Data. The second area visualizes the model’s pipeline to encode diverse sources and illustrates how LLMs are applied for data analysis. The third area describes how the model consolidates outputs from both embeddings and LLM analysis for use in subsequent stages. The model will perform multi-task learning: our RiskLabs will predict the Volatility of different terms and VaR in the meantime.
2.1 Multi-Source Data Input
Previous research has shown that textual information in a company’s financial statements can be used to predict the level of risk in its stock Kogan et al. (2009). In the current landscape, CEOs communicate not only through written mediums like press releases and financial reports but also through investor meetings and earnings conference calls, often captured in audio and video formats. Technological advancements have made it possible to archive and access audio recordings of investor meetings and videos of CEOs’ speeches, available to both organizations and individuals. Additionally, research in natural language processing (NLP) indicates that a CEO’s affective state, tone, and mood, as conveyed through voice cues, can complement verbal information and enhance risk prediction Qin and Yang (2019). Therefore, relevant earnings conference call data form a vital component of our dataset. In detail, each earnings conference call data sample consists of two parts: an audio record and a corresponding transcript.
Furthermore, the literature suggests that combining historical numerical data can enhance the accuracy of financial predictions, considering the varied frequencies of multiple data sources Sawhney et al. (2020). To accommodate this, we gather historical price data for the 30 days preceding earnings announcements, acknowledging the different frequencies of data sources like daily price updates and quarterly earnings notices. This historical price data represents our second key data source.
Additionally, prior research has often overlooked the influence of earnings conference call releases on long-term forecasts. While earnings conference calls can cause significant short-term volatility in financial risk, their quarterly frequency diminishes their long-term forecast impact. To address this, we also compile daily relevant news texts centered around the firm, serving as our third crucial data source. This approach aims to provide a more comprehensive view of the company’s financial standing, considering both short-term events and long-term trends.
In this study, the dataset is from the publicly available S&P 500 earnings conference call dataset as constructed by Qin and Yang (2019). This dataset comprises both audio recordings and their corresponding text transcripts from the 2017 earnings conference calls of 500 major public companies listed in the S&P 500 Index and traded on U.S. stock exchanges. Originally, the dataset contained 2,243 conference calls from 2017. However, due to the challenge of noisy and imprecise alignments between audio and text, a significant portion of this raw data was deemed unusable. Consequently, we focused on 576 unique instances where the audio recordings were accurately and closely aligned with the text transcripts. These instances represent profitable conference calls and encompass a total of 88,829 sentences where the text and audio data were well-matched. To enrich this dataset, we collected corresponding time-series data and conducted thorough research to identify relevant news items pertaining to these companies and the dates of their ECC releases. This extensive process of data collation and curation led to the construction of the comprehensive dataset employed in our experiments.
2.2 Response Variables
RiskLabs is a multi-task training model that trains and predicts the volatility of different terms and VaR. In the model, our first task is to predict the four different terms of volatility, which are 3-day, 7-day, 15-day, and 30-day volatility. We focused on different volatility terms for the following reasons: (1) Multiple terms could reflect the level of stability in different time horizons. In our mind, we consider that 3-day, 7-day could reflect the short-term risk, 15-days could reflect the mid-term Risk, and 30-days could reflect the long-term risk; (2) 30-day volatility will link to the VIX; (3) Volatility will be an essential part for the portfolio management, specifically in portfolio weight adjustment and asset selection.
$x=F^{-1}(p)$
The $F(\cdot)$ is the cumulative loss distribution, $p$ is the percentile we set, and $x$ is the VaR. From the idea of Quantile Regression, we can have:
Calculating and estimating VaR can help the company better deal with financial risks and avoid extreme scenarios in the future.
2.3 Earnings Conference Call Encoder
The Earnings Conference Call Encoder module is composed of four key components: 1) Audio Encoding; 2) Transcript Encoding; 3) Earnings Conference Call Analyzer; and 4) Additive Multi-modal Fusion. In the first two components, audio and textual data are initially converted into vector space representations using an appropriate pre-trained model. Subsequently, a multi-head attention mechanism is employed to extract salient features. The third component involves the utilization of a large language model to summarize and abstract the text. This process aims to distill the essential information contained in the text and transform it into a suitable vector format. Lastly, the Additive Multi-modal Fusion integrates these feature sets, effectively combining the information extracted from both audio and textual data. This holistic approach ensures a comprehensive analysis of the earnings conference call, leveraging multiple data modalities for a more robust encoding.
2.3.1 Audio Encoding
The Audio Encoding component consists of a pre-trained audio embedding model, a Multi-Head Self-Attention, and an Average Pooling layer. Firstly, we extracted audio embedding using the Wav2vec2, a transformer-based Large Language Model proposed by Baevski et al. (2020). The Wav2vec2 model has been recognized for its effectiveness, showing that learning from speech audio alone can surpass some of the best semi-supervised methods in performance while maintaining conceptual simplicity. Wav2vec2 operates by masking the speech input in the latent space and engaging in a contrastive task based on a quantization of these latent representations, which are learned concurrently. Following this, the audio embeddings are processed through the Multi-Head Self-Attention mechanism. This step is crucial for distilling specific audio features, which are then poised for integration with features from other data modalities, ensuring a comprehensive and nuanced analysis.

### v2

# RiskLabs: Predicting Financial Risk Using Large Language Model based on Multimodal and Multi-Sources Data

@@ Abstract.
The integration of Artificial Intelligence (AI) techniques, particularly large language models (LLMs), in finance has garnered increasing academic attention. Despite progress, existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&\&&A), and stock movement prediction (binary classification), the application of LLMs to financial risk prediction remains underexplored. Addressing this gap, in this paper, we introduce RiskLabs, a novel framework that leverages LLMs to analyze and predict financial risks. RiskLabs uniquely integrates multimodal financial data—including textual and vocal information from Earnings Conference Calls (ECCs), market-related time series data, and contextual news data—to improve financial risk prediction. Empirical results demonstrate RiskLabs’ effectiveness in forecasting both market volatility and variance. Through comparative experiments, we examine the contributions of different data sources to financial risk assessment and highlight the crucial role of LLMs in this process. We also discuss the challenges associated with using LLMs for financial risk prediction and explore the potential of combining them with multimodal data for this purpose.  [^1] Position Paper [^2] Present at International Workshop on Multimodal Financial Foundation Models (MFFMs) @ ICAIF’24333https://sites.google.com/view/iwmffm2024/accepted-papers?authuser=1444 This position work was completed in April 2024.
Large Language Model, Financial Risk Forecasting, Multimodal Learning, Multi-source Data
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance; November 15, 2024; Brooklyn, NY,isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The integration of artificial intelligence (AI) technology for financial risk prediction has been a longstanding focus of both industry and academia. Previously, researchers used machine learning techniques to analyze historical stock price data for risk prediction, such as using SVM to predict credit ratings based on financial statement ratios (Khaidem et al., 2016; Lee, 2007) and using tree-based models to explore the correlation between economic indicators and the performance of various asset classes (Gu et al., 2020). Recently, researchers have increasingly extracted trading signals from publicly available information for risk prediction. For example, financial reports and media news are used to analyze sentiment in news articles, capture market expectations for stock movements (Souma et al., 2019; Mohan et al., 2019; Liapis et al., 2023), risk factor analysis (Zhang, 2020), and identify early warning signals from large amounts of textual data (Ahbali et al., 2022; Wang et al., 2023). With the rise of multimodal learning, more unstructured multimedia data, such as audio recordings of earning conference calls (Qin and Yang, 2019; Yang et al., 2020) are utilized for stock volatility prediction.
While supervised learning methods have proven effective, they are highly task-specific and lack adaptability (Singh et al., 2016), with their predictive performance constrained by the amount of input data and the size of parameters they can handle. However, the emergence of large language models (LLMs) represents a paradigm shift in overcoming these challenges. Equipped with vast knowledge bases and advanced zero-shot learning capabilities, LLMs can perform a wide range of text-related tasks, such as summarization (Zhang et al., 2024a), question answering (Wei et al., 2022), and sentiment analysis (Zhang et al., 2023), without requiring task-specific training. In the financial domain, LLMs have been widely utilized for tasks such as financial text analysis (Yang et al., 2023) and financial report generation (Xie et al., 2024). Furthermore, LLM-based agent systems have been employed for stock trading (Yu et al., 2023; Zhang et al., 2024b, c). These studies highlight the versatility and impact of LLMs in finance, emphasizing their growing significance in enhancing business operations and decision-making processes.
@@ 1. Introduction
Figure 1. This figure illustrates the RiskLabs Framework. The model accepts multiple-source inputs: Earnings Conference Call Audio & Transcript, Daily News, and Time Series Data. The second area visualizes the model’s pipeline to encode diverse sources and illustrates how LLMs are applied for data analysis. The third area describes how the model consolidates outputs from both embeddings and LLM analysis for use in subsequent stages. The model will perform multi-task learning: our RiskLabs will predict the Volatility of different terms and VaR in the meantime.
Despite the expanding body of work on LLM applications in finance, there remains a significant gap in exploring their potential for financial risk prediction, particularly for metrics like stock volatility and Value at Risk (VaR). In addition, since LLMs are generative models, the challenge, of using them for precise numerical regression (Song et al., 2024), also persists in financial risk prediction tasks. Therefore, our research aims to fill this gap by leveraging LLMs for financial risk prediction, contributing valuable insights to understanding market dynamics and ensuring sustainable success across various sectors. This position paper is interested in exploring the following research questions (RQs):
  • RQ1: What role do large language models play in financial risk prediction?
  • RQ2: How does the predictive performance of Large Language Models (LLMs) compare with other artificial intelligence techniques in forecasting risk metrics?
  • RQ3: How can multiple inputs of varying data types be effectively incorporated and balanced?
To address the issues previously mentioned, we propose the RiskLabs framework, designed to comprehensively understand the complexities of the investment environment. This framework ingests and analyzes multimodal data from multiple sources, providing a holistic view of the factors influencing financial markets. The multimodal data sources include: (1) earnings conference call transcripts, offering insights into corporate performance and future prospects; (2) audio from these calls, capturing tone and sentiment not evident in the text; (3) time series data for historical and real-time market trend analysis and (4) media news. Leveraging LLMs, RiskLabs extracts essential information from each modality and applies multimodal fusion techniques to integrate these features, enhancing financial risk prediction.
RiskLabs is equipped with four key modules to effectively process these diverse data streams: a) The Earnings Conference Call Encoder: leveraging LLM to handle data related to earnings conference call (1 and 2); b) The News-Market Reactions Encoder: establishing a pipeline through LLM to collect and interpret news data; c) The Time-Series Encoder: organizing and analyzing time-related data; and d) The Multi-Task Prediction: which amalgamates outputs from the aforementioned modules for multifaceted prediction. This synthesis of varied inputs allows the model to provide a nuanced understanding of the investment landscape, blending quantitative market data with qualitative insights. Such an integrated methodology is crucial for precise forecasting and risk assessment, a key to navigating the often turbulent and complex terrain of financial markets. By incorporating a diverse range of inputs and using processing steps, RiskLabs solves a multitasking problem: forecasting volatility over various intervals (3 days, 7 days, 15 days, and 30 days), as well as the value at risk (VaR).
Our contributions are encapsulated as follows: 1) We utilize LLM to develop the RiskLabs framework, thereby addressing the shortfall in applying LLMs to financial risk prediction. 2) The versatility of our model lies in its capability to seamlessly integrate a variety of financial multimodal data from multiple sources, consequently bolstering risk prediction accuracy. 3) The efficacy of our framework is evidenced by experimental results, demonstrating its effectiveness in forecasting financial risks. 4) Furthermore, we provide a comprehensive analysis of the pivotal role LLM should assume in the realm of financial risk prediction.
2. RiskLabs Framework
Figure 1 illustrates the RiskLabs Framework, designed to handle multiple data types surrounding the financial information including audio, text, and time-series from different sources. The framework comprises four main modules: 1) Earnings Conference Call Encoder; 2) Time-Series Encoder; 3) Relevant News Encoder; and 4) Multimodal Fusion and Multi-Task Prediction. This section outlines the format of the data input and details each module. In brief, the Earnings Conference Call Encoder, Time-Series Data Encoder, and Relevant News Encoder are utilized to extract features from various data types. These features are fused data that undergoes processing and modeling, after which it is fed into the Multi-Task Prediction Block. This block is responsible for forecasting both volatility for different intervals and VaR (Value at Risk) values.
Figure 2. visualizes the process of fine-grained information extraction from ECC transcript.
2.1. Earnings Conference Call Encoder
The Earnings Conference Call Encoder module consists of three main components: Audio Encoding, Transcript Encoding, and Earnings Conference Call Analyzer.
2.1.1. Audio Encoding